#52 A simple ohmmeter is made by connecting a 1.50V flash- light battery in series with a resistance R and an ammeter that

media%2F8c6%2F8c63afdb-55f5-4c27-853f-d7

please complete each section of the problem:a,b,c,d
Answer #1

When the clips are shorted together

V=IR

\Rightarrow 1.5=1\times 10^{-3}\times R

1.5 1 × 10-3 = 1500 Ω

So, the value of R is 1500 Ohm.

PART A: When Current is 10% of its maximum value of 1.5 mA.

V = 1(R + r)

1 x 103 x 10 100 1.5 = (1500+r)

1500+r-15000

Y-15000-1500-13500 Ω

So, for 10% of max current, the resistance should be 13500 Ohm.

PART B: When Current is 50% of its maximum value of 1.5 mA.

V = 1(R + r)

1 x 10-3 × 50 100 1.5 = (1500+r)

1500+r-3000

Y-3000-1500-1500 Ω

So, for 50% of max current, the resistance should be 1500 Ohm.

PART C: When Current is 90% of its maximum value of 1.5 mA.

V = 1(R + r)

1.5=\frac{1\times 10^{-3}\times 90}{100}(1500+r)

\Rightarrow 1500+r=1667

\Rightarrow r=1667-1500=167\ \Omega

So, for 90% of max current, the resistance should be 167 Ohm.

PART D:

If the ammeter has an internal resistance of 20 ohms. Then

V = 1(R + r)

\Rightarrow 1.5=1\times 10^{-3}\times (R+20)

\Rightarrow R=1500-20=1480\ \Omega

So, if the ammeter has an internal resistance of 20 Ohms, the value of R would be 1480 ohms.